A projectile is launched at a speed of 20 m/s and at an angle of 25 degrees above the horizontal. What is the maximum height of the projectile?


1
Expert's answer
2021-10-20T16:39:55-0400

The maximum height is


"h=\frac{v^2\sin^2\theta}{2g}=3.65\text{ m}."